OpenStudy (anonymous):

Evergreen landscaping bought 8 tons of topsoil, 5 tons of mulch, and 2 tons of pea gravel for $2943. The next week the firm bought 5 tons of topsoil, 4 tons of mulch, and 5 tons of pea gravel for $2700. Pea gravel costs$26 less per ton than topsoil. Find the cost per ton for each item.

OpenStudy (texaschic101):

g = t - 26 8t + 5m + 2g = 2943 8t + 5m + 2(t - 26) = 2943 8t + 5m + 2t - 52 = 2943 10t + 5m = 2943 + 52 10t + 5m = 2995 5t + 4m + 5g = 2700 5t + 4m + 5(t - 26) = 2700 5t + 4m + 5t - 130 = 2700 10t + 4m = 2700 + 130 10t + 4m = 2830 10t + 5m = 2995 10t + 4m = 2830 -->(-1) ------------------ 10t + 5m = 2995 -10t - 4m = -2830 (result of multiplying by -1) -----------------add m = 165 10t + 5m = 2995 10t + 5(165) = 2995 10t + 825 = 2995 10t = 2995 - 825 10t = 2170 t = 217 g = t - 26 g = 217 - 26 g = 191 check.... 8t + 5m + 2g = 2943 8(217) + 5(165) + 2(191) = 2943 1736 + 825 + 382 = 2943 2943 = 2943 (correct) topsoil (t) costs $217 per ton mulch (m) costs $165 per ton pea gravel (g) costs $191 per ton

OpenStudy (texaschic101):

any questions ?